OpenStudy (anonymous):

identify the point of intersection, using one of the following methods: Elimination, subsitiution, graphing X starts at 0 and goes up by 10 each rep for up to 15 reps To find y for the first its .4(X)+50 To find y for the second line its 4.4(X)+0

myininaya (myininaya):

what is rep?

OpenStudy (anonymous):

repetition so it goes up by 10 15 times

myininaya (myininaya):

so are you saying the equations are: y=.4x+50 and y=4.4x ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok yes so i don't have to read that rep stuff lol

OpenStudy (anonymous):

idk but its within 8-12 bottles

OpenStudy (anonymous):

yes its 10 bottles bottles are the item we are making a graph on and x = bottles

OpenStudy (anonymous):

no its not ten =0

myininaya (myininaya):

just set 4.4x=.4x+50 4.4x-.4x=50 4x=50 x=50/4 x=25/2 so y=4.4x=4.4(25/2)=44/10(25/2)=22/5(25/2)=5(11)=55 so they intersect at (25/2,55)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

hey derp u just saw my question bt i asked u something can u please answer me please

myininaya (myininaya):

to find where they intersect all i did was set the two lines equal to each other i want to find the x when the two lines are the same (so the equal sign means that) and i solved for x and then it didn;t matter which line i used after that to find y both lines would give me the same y

myininaya (myininaya):

yes i just replaced that one y with y=4.4x substitution

OpenStudy (anonymous):

would i get full credit for expaining work also what is the function of the two

myininaya (myininaya):

in the directions about it said to find the intersection it is (25/2,55) i don't know what you mean by the function of the two?

OpenStudy (anonymous):

ok its function for 1st line and function for 2 and if the points matter two of them from each are (1) 10,52 20,54 (2) 10,44 20, 88

OpenStudy (anonymous):

hey derp answer me please

OpenStudy (anonymous):

ok while hes looking at mine ill help u

myininaya (myininaya):

wait are you saying for the first line the some of the ordered pairs are {(10,52),(20,54)} 2nd line {(10,44),{20,88)} ?

OpenStudy (anonymous):

by the scale they gave us these are for each 10 bottles but for the second part which i started the question it is ok because plugin them into a graph got around what you said.

OpenStudy (anonymous):

we were told to go by 10

myininaya (myininaya):

the above are some of the ordered pairs on the line right? you are trying to get me to check your lines, correct?

OpenStudy (anonymous):

the function is what im clueless about and i think it has some relation with that right?

myininaya (myininaya):

are the following ordered pairs on a line: (10,52) and (20,54) are the following ordered pairs on the other line: (10,44) and (20,88)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So?

myininaya (myininaya):

ok so since we know two points on the first line, the slope is \[\frac{54-52}{20-10}=\frac{2}{10}=\frac{1}{5}\] so we have y=1/5*x+b we need to know the y-intercept which we are calling b in the line before we know a point on the line (10,52) x=10 and y=52 so 52=1/5*10+b we can solve for b 52=2+b 50=b so the equation of line 1 is y=1/5*x+50 now second line we know two points on it as well so we can find the slope the slope is \[\frac{88-44}{20-10}=\frac{44}{10}=\frac{22}{5}\] so we have y=22/5*x+b we know a point on this line so we can find the y-intercept which we called b in the line before x=10 y=44 44=22/5*10+b so we can solve for b 44=22(2)+b 44=44+b b=0 so the equation of the line is y=22/5*x

myininaya (myininaya):

so we have Line 1: y=1/5*x+50 and Line 2: y=22/5*x 22/5*x=1/5*x+50 (by substitution) 22/5*x-1/5*x=50 21/5*x=50 x=50*5/21 x=250/21

myininaya (myininaya):

so y=22/5*250/21=1100/21

OpenStudy (anonymous):

so the 25/2,55 changed?

myininaya (myininaya):

so the point of intersection is (250/21,1100/21)

myininaya (myininaya):

yes we have different lines so the point of intersection will be different

OpenStudy (anonymous):

are you sure its not closer because on the graph i made its near the start

myininaya (myininaya):

? i found the lines by the ordered pairs you gave me those ordered pairs lie on those two lines right with (10,52) and (20,54) lying on one line and with (10,44) and (20,88) lying on the other?

OpenStudy (anonymous):

yea but 10 is 10 of x you can plug 1,2,3,- 150